OpenStudy (anonymous):

2/3y+6=y can someone help me ?!?

OpenStudy (anonymous):

2/3y+6=y 3y(2/3y +6)=3y(y) 2+18y=4y subtract -18y to both sides 2=-14y divide -14 to both sides y= -1/7

OpenStudy (anonymous):

-18 , 18 , 1/3 and -1/3 are my only options doe